# boxedasideAside

### v1

@@ A.3 Deformation theory of $\overline{M}_{E,g;n}$
    The map $\phi:X\to M_{E,g}$: By definition, this is the same as specifying a stable map $(\mathcal{C}\to X,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ such that $f$ is degree 1, all the automorphism groups of $\mathcal{C}$ above $\Sigma$ have order 2, and $\mathcal{C}\to X$ is smooth. There is a universal section $s:E\times X\to\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$, where $\mathcal{L}|_{E\times X}$ is the pullback of $\mathcal{L}$ along the map $E\times X\to E\times\textup{Pic}^{g-1}(E)$. We define $D\subset\mathcal{L}|_{E\times X}$ as the preimage of $s(E\times X)\subset\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$ along the squaring map $\mathcal{L}|_{E\times X}\to\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$. We then let $\mathcal{C}\coloneqq[D/(\mathbb{Z}/2\mathbb{Z})]$, where $\mathbb{Z}/2\mathbb{Z}$ acts on $D\subset\mathcal{L}|_{E\times X}$ by multiplication by ±$\pm 1$. The projection $D\to E\times X$ exhibits $E\times X$ as the coarse quotient $C=D/(\mathbb{Z}/2\mathbb{Z})$. Let $\Sigma\subset E\times X$ be the branch divisor of the double cover $D\to C$, which is the same as the vanishing locus of the universal section $s$. Because of the requirement that sections $r$ corresponding to points $(\mathcal{M},r)\in X$ have reduced vanishing loci, $\Sigma\to X$ is étale. Checking that $(\mathcal{C}\to C\to X,\Sigma)$ is a smooth twisted $(g-1)$-pointed curve can be done using the description of $\mathcal{C}$ as the quotient $[D/(\mathbb{Z}/2\mathbb{Z})]$ and the fact that $\Sigma\to X$ is étale, as these facts ensure that we have the local description we want for Definition A.1.
    We define $f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z}$ by letting $\mathcal{C}\to E$ be the projection $\mathcal{C}\to C=E\times X\to E$ and letting $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$ classify the étale double cover $D\to\mathcal{C}$. The fact that $f$ is representable follows from the fact that $D$ is a scheme. The map on coarse spaces $C\to E$ is just the projection $E\times X\to E$, so it is a stable map of degree 1. The last thing we have to check is that the points in $\mathcal{C}$ above $\Sigma$ have stabilizer $\mathbb{Z}/2\mathbb{Z}$. This is true because the preimage of $\Sigma$ in $D$ is precisely the intersection of $D$ with the zero section of $\mathcal{L}|_{E\times X}$, which is stabilized by the $\mathbb{Z}/2\mathbb{Z}$-action on $D$. We thus have a twisted stable map $(\mathcal{C}\to X,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})\in M_{E,g}$, defining a map $\phi:X\to M_{E,g}$.
@@
    $\phi$ is surjective: This follows from the well-known fact that double covers of a curve can all be constructed as the square root of a section of a line bundle as above, which we will explain briefly below. In the following, we reset the notation we used earlier in the proof. Let $(\mathcal{C}\to{\rm Spec}k,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be any point of $M_{E,g}$. Then $\mathcal{C}$ has coarse space identified with $E$ by $f$, and $\mathcal{C}=[D/(\mathbb{Z}/2\mathbb{Z})]$, where $D\to\mathcal{C}\to E$ is the double cover classified by the map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$. We want to show that $h:D\to E$ can be produced using the square root construction above. Let $\mathcal{M}\coloneqq(h_{*}\mathcal{O}_{D}/\mathcal{O}_{E})^{\vee}$. Because 2 is invertible, the trace map provides a splitting $h_{*}\mathcal{O}_{D}\to\mathcal{O}_{E}$ of the exact sequence $0\to\mathcal{O}_{E}\to h_{*}\mathcal{O}_{D}\to\mathcal{M}^{\vee}\to 0$, so that we get a section $\mathcal{M}^{\vee}\to h_{*}\mathcal{O}_{D}$. This extends to a surjection of $\mathcal{O}_{E}$-algebras ${\rm Sym}_{\mathcal{O}_{E}}^{\bullet}(\mathcal{M}^{\vee})\to h_{*}\mathcal{O}_{D}$, which provides a closed embedding $D\xhookrightarrow{}\mathcal{M}$. The ideal $\ker({\rm Sym}_{\mathcal{O}_{E}}^{\bullet}(\mathcal{M}^{\vee})\to h_{*}\mathcal{O}_{D})$ is preserved by the $\mathbb{Z}/2\mathbb{Z}$-action acting by $-1$ on $\mathcal{M}^{\vee}$, so $D$ is preserved by the $\mathbb{Z}/2\mathbb{Z}$-action. This means that its image under the squaring map $\mathcal{M}\to\mathcal{M}^{\otimes 2}$ maps by degree 1 to $E$ and is thus isomorphic to $E$. Thus, we have a section $r:E\to\mathcal{M}^{\otimes 2}$ whose square root gives us the double cover $D\to E$. Because $D$ is smooth, the vanishing locus of $r$ must be reduced, and we know that a point $(\mathcal{M},r)\in X$ maps to our point $(\mathcal{C}\to{\rm Spec}(k),\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ of $M_{E,g}$. We conclude that $h$ is surjective and that $M_{E,g}$ is irreducible.
@@
We first show that $M_{E,g;n}$ is irreducible. Since it is smooth, it suffices to show that it is connected.
@@
We consider the forgetful morphism $\pi:M_{E,g;n}\to M_{E,g}$. The fiber is the configuration space which is connected. If we let $\Delta$ denote the weak diagonal in $E^{n}$, and let $R$ be the ramification locus of the fixed double cover, then the fiber over the double cover is $((E-R)^{n}-\Delta_{weak})/S_{n}$, which is connected. Also, the image of $\pi$ is $M_{E,g}$, which is connected. Therefore, $M_{E,g;n}$ is connected, hence irreducible. By dimension, since $M_{E,g;n}$ is dense in $\overline{M}_{E,g;n}$, it follows that $\overline{M}_{E,g;n}$ is connected, hence irreducible. ∎

### v2

@@ A.3 Deformation theory of $\overline{M}_{E,g;n}$
    The map $\phi:X\to M_{E,g}$: By definition, the map $\phi$ can be defined by specifying a stable map $(\mathcal{C}\to X,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ such that $f$ is degree 1, all the automorphism groups of $\mathcal{C}$ above $\Sigma$ have order 2, and $\mathcal{C}\to X$ is smooth. There exists a universal section $s:E\times X\to\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$, where $\mathcal{L}|_{E\times X}$ is the pullback of $\mathcal{L}$ along the map $E\times X\to E\times\textup{Pic}^{g-1}(E)$. We define $D\subset\mathcal{L}|_{E\times X}$ as the preimage of $s(E\times X)\subset\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$ along the squaring map $\mathcal{L}|_{E\times X}\to\left(\mathcal{L}|_{E\times X}\right)^{\otimes 2}$. We then let $\mathcal{C}\coloneqq[D/(\mathbb{Z}/2\mathbb{Z})]$, where $\mathbb{Z}/2\mathbb{Z}$ acts on $D\subset\mathcal{L}|_{E\times X}$ by multiplication by $\pm 1$. The projection $D\to E\times X$ exhibits $E\times X$ as the coarse quotient $C=D/(\mathbb{Z}/2\mathbb{Z})$. Let $\Sigma\subset E\times X$ be the branch divisor of the double cover $D\to C$, which is the same as the vanishing locus of the universal section $s$. Because of the requirement that sections $r$ corresponding to points $(\mathcal{M},r)\in X$ have reduced vanishing loci, $\Sigma\to X$ is étale. Checking that $(\mathcal{C}\to C\to X,\Sigma)$ is a smooth twisted (g1)(g-1)-pointed curve can be done using the description of $\mathcal{C}$ as the quotient $D/(\mathbb{Z}/2\mathbb{Z})]$ and the fact that $\Sigma\to X$ is étale, as these facts ensure that we have the local description we want for Definition A.1.
    We define $f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z}$ by letting $\mathcal{C}\to E$ be the projection $\mathcal{C}\to C=E\times X\to E$ and letting $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$ classify the étale double cover $D\to\mathcal{C}$. The fact that $f$ is representable follows from the fact that DD is a scheme. The map on coarse spaces $C\to E$ is just the projection $E\times X\to E$, so it is a stable map of degree 1. The last thing we have to check is that the points in $\mathcal{C}$ above $\Sigma$ have stabilizer $\mathbb{Z}/2\mathbb{Z}$. This is true because the preimage of $\Sigma$ in DD is precisely the intersection of $D$ with the zero section of $\mathcal{L}|_{E\times X}$, which is stabilized by the $\mathbb{Z}/2\mathbb{Z}$-action on $D$. We thus have a twisted stable map $(\mathcal{C}\to X,\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})\in M_{E,g}$, defining a map $\phi:X\to M_{E,g}$.
@@
    $\phi$ is surjective: This follows from the well-known fact that double covers of a curve can all be constructed as the square root of a section of a line bundle as above, which we will explain briefly below. In the following, we reset the notation we used earlier in the proof. Let $(\mathcal{C}\to{\rm Spec}(k),\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ be any point of $M_{E,g}$. Then $\mathcal{C}$ has coarse space identified with $E$ by $f$, and $\mathcal{C}=[D/(\mathbb{Z}/2\mathbb{Z})]$, where $D\to\mathcal{C}\to E$ is the double cover classified by the map $\mathcal{C}\to B\mathbb{Z}/2\mathbb{Z}$. We want to show that $h:D\to E$ can be produced using the square root construction above. Let $\mathcal{M}\coloneqq(h_{*}\mathcal{O}_{D}/\mathcal{O}_{E})^{\vee}$. Because 2 is invertible, the trace map provides a splitting $h_{*}\mathcal{O}_{D}\to\mathcal{O}_{E}$ of the exact sequence $0\to\mathcal{O}_{E}\to h_{*}\mathcal{O}_{D}\to\mathcal{M}^{\vee}\to 0$, so that we get a section $\mathcal{M}^{\vee}\to h_{*}\mathcal{O}_{D}$. This extends to a surjection of $\mathcal{O}_{E}$-algebras ${\rm Sym}_{\mathcal{O}_{E}}^{\bullet}(\mathcal{M}^{\vee})\to h_{*}\mathcal{O}_{D}$, which provides a closed embedding $D\xhookrightarrow{}\mathcal{M}$. The ideal $\ker\left({\rm Sym}_{\mathcal{O}_{E}}^{\bullet}(\mathcal{M}^{\vee})\to h_{*}\mathcal{O}_{D}\right)$ is preserved by the $\mathbb{Z}/2\mathbb{Z}$-action acting by 1-1 on $\mathcal{M}^{\vee}$, so DD is preserved by the $\mathbb{Z}/2\mathbb{Z}$-action. This means that its image under the squaring map $\mathcal{M}\to\mathcal{M}^{\otimes 2}$ maps by degree 1 to $E$ and is thus isomorphic to $E$. Thus, we have a section $r:E\to\mathcal{M}^{\otimes 2}$ whose square root gives us the double cover $D\to E$. Because DD is smooth, the vanishing locus of $r$ must be reduced, and we know that a point $(\mathcal{M},r)\in X$ maps to our point $(\mathcal{C}\to{\rm Spec}(k),\Sigma,f:\mathcal{C}\to E\times B\mathbb{Z}/2\mathbb{Z})$ of $M_{E,g}$. We conclude that $h$ is surjective and that $M_{E,g}$ is irreducible.
@@
We show that $M_{E,g;n}$ is connected. Since it is smooth, this implies that it is irreducible.
@@
We consider the forgetful morphism $\pi:M_{E,g;n}\to M_{E,g}$. The fiber is the configuration space, which is connected. If we let $\Delta$ denote the weak diagonal in $E^{n}$, and let $R$ be the branch locus of the fixed double cover, then the fiber over the double cover is $((E-R)^{n}-\Delta)/S_{n}$, which is connected. Also, the image of $\pi$ is $M_{E,g}$, which is connected. Therefore, $M_{E,g;n}$ is connected, hence irreducible. By dimension, since $M_{E,g;n}$ is dense in $\overline{M}_{E,g;n}$, it follows that $\overline{M}_{E,g;n}$ is connected, hence irreducible. ∎
Proposition A.15.
We have the following morphism:
$\overline{M}_{E,g}^{{\rm red}}\xrightarrow{\varphi}\overline{\mathcal{M}}_{g,2g-2}\xrightarrow{\pi}\overline{\mathcal{M}}_{g}$
The first map $\varphi$ sends a cover $D\to C$ to its source curve $D$, with the 2g22g-2 ramification points as marked points. The second map $\pi$ forgets the marked points and stabilizes the curve. Then $\pi\circ\varphi$ is quasi-finite on the smooth locus of $\overline{M}^{{\rm red}}_{E,g}$ and generically finite on each boundary component of $\overline{M}^{{\rm red}}_{E,g}$.
Proof.
Observe that $\pi\circ\varphi$ factors through $\overline{M}_{E,g}/{\rm Aut}(E)]$, and we already know that the map from $\overline{M}_{E,g}^{{\rm red}}$ to $\overline{M}_{E,g}/{\rm Aut}(E)]$ is finite. Therefore, it is enough to show the statement for the map from $\overline{M}_{E,g}/{\rm Aut}(E)]$ to $\overline{\mathcal{M}}_{g}$. Hence, we want to show that the forgetful map has finite fibers on the smooth locus and generically on each boundary component. Let $g\geq 2$.
  1. 1.
    We first show quasifiniteness on the smooth locus. We know that $M_{E,g}/{\rm Aut}(E)]$ parametrizes double covers $D\xrightarrow{\eta}E$ such that $D$ has genus $g$. Moreover, $D_{1}\xrightarrow{\eta_{1}}E$ and $D_{2}\xrightarrow{\eta_{2}}E$ correspond to the same point of $M_{E,g}/{\rm Aut}(E)]$ if there are isomorphisms $f_{1}:D_{1}\to D_{2}$ and $f_{2}:E\to E$ such that $\eta_{1}=f_{2}\circ\eta_{2}\circ f_{1}$.
    Consider a curve $D$ in the image under $\pi\circ\varphi$ of the smooth locus of $\overline{M}^{{\rm red}}_{E,g}$, and let $\eta:D\to E$ be the data of one of the preimages. We will show that it has only finitely many preimages in $M_{E,g}/{\rm Aut}(E)]$. First, note that each double cover $\eta:D\to E$ gives us an element $\sigma\in{\rm Aut}(D)$ which is of order $2$. If we let $\eta^{*}$ and $\sigma^{*}$ to be the induced map on the function field $K(E)$ and $K(D)$ respectively, then we have that $\eta^{*}(K(E))=K(D)^{\sigma^{*}}$. Since $g\geq 2$, we know that ${\rm Aut}(D)$ is a finite group. Therefore, there are finitely many choices for $\sigma$ and hence finitely many choices for $\eta^{*}(K(E))$.
    Now, we show that $\eta^{*}(K(E))$ determines the preimage. Suppose, we have two preimages $D\xrightarrow{\eta_{0}}E$ and $D\xrightarrow{\eta_{1}}E$ that satisfy $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$. Then there is $\tau\in{\rm Aut}(K(E))$, for which $\eta_{1}^{*}\circ\tau=\eta_{0}^{*}$ (this $\tau$ is obtained from the composition $K(E)\xrightarrow{\eta_{0}^{*}}\eta_{0}^{*}(K(D))\xrightarrow{{\eta_{1}^{*}}^{-1}}K(E)$, which is well defined since $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$). Next, there is $f\in{\rm Aut}(E)$ for which $f^{*}=\tau$, so $f\circ\eta_{1}=\eta_{0}$. This shows that $D\xrightarrow{\eta_{0}}E$ and $D\xrightarrow{\eta_{1}}E$ correspond to the same point of $M_{E,g}/{\rm Aut}(E)]$. We conclude that the forgetful map from $M_{E,g}/{\rm Aut}(E)]$ to $\overline{\mathcal{M}}_{g}$ is quasi-finite on the smooth locus.
  2. 2.
    Next, consider a boundary component, either $\Xi_{g_{1},g_{2}-1}$ or $\Delta_{g_{1},g_{2}}$. Curves in the generic part of any boundary component are the clutching of two smooth curves, except for one component of $\Xi_{1,g-2}$, when it is the clutching of three smooth curves. Consider a curve $D\in\overline{\mathcal{M}}_{g}$ coming from a generic point of a boundary component. Let $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ be a preimage of $D$, where $C_{0}$ is the clutching of $E$ with $\mathbb{P}^{1}$ (say $P_{0}$ of $E$ with $Q_{0}$ of $\mathbb{P}^{1}$), and DD is the stabilization of $D^{\prime}$ after forgetting the ramification points. Let $D_{a}=\eta_{0}^{-1}(E)$ and $D_{b}=\eta_{0}^{-1}(\mathbb{P}^{1})$, so $g(D_{a})=g_{1}$ and $g(D_{b})=g_{2}$. Then $D^{\prime}$ is the clutching of $D_{a}$ and $D_{b}$, when $\eta_{0}^{-1}(P_{0})$ is identified with $\eta_{0}^{-1}(Q_{0})$. Notice that if $g_{2}\geq 1$, then $D=D^{\prime}$. On the other hand, if $g_{2}=0$, then it must be the case that $(D^{\prime}\to C_{0})\in\Xi_{g-1,0}$. In this case, DD is $D_{a}$ with two points in $\eta_{0}^{-1}(P_{0})$ identified.
    1. (a)
      If $g_{1},g_{2}\geq 2$. Then, since a curve of genus $\geq 2$ has a finite automorphism group, we have only finitely many options for $\sigma_{a}$ and $\sigma_{b}$. We show that $\sigma_{a},\sigma_{b}$ uniquely determine the map $D^{\prime}\to C_{0}\to E$.
      Let $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ be another preimage of $D$ under $\pi\circ\varphi$. Suppose that we have $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$ and $\eta_{0}^{*}(K(\mathbb{P}^{1}))=\eta_{1}^{*}(K(\mathbb{P}^{1}))$. Then, there exist $h_{a}\in{\rm Aut}(E)$ and $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{1}^{*}\circ h_{a}^{*}=\eta_{0}^{*}$ and $\eta_{1}^{*}\circ h_{b}^{*}=\eta_{0}^{*}$, so $h_{a}\circ\eta_{1}=\eta_{0}|_{D_{a}}$ and $h_{b}\circ\eta_{1}=\eta_{0}|_{D_{b}}$. Recall that $D^{\prime}$ is the clutching of $D_{a}$ and $D_{b}$, when $\eta_{0}^{-1}(P_{0})\subseteq D_{a}$ is identified with $\eta_{0}^{-1}(Q_{0})\subseteq D_{b}$. Since $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage of $D$, we see that DD is also the clutching of $D_{a}$ and $D_{b}$, when $\eta_{1}^{-1}(P_{1})$ is identified with $\eta_{1}^{-1}(Q_{1})$. This means that $\eta_{0}^{-1}(P_{0})=\eta_{1}^{-1}(P_{1})$ and $\eta_{0}^{-1}(Q_{0})=\eta_{1}^{-1}(Q_{1})$. We see that $h_{a}(P_{1})=P_{0}$ and $h_{b}(Q_{1})=Q_{0}$. Thus the map $h=h_{a}\cup h_{b}$ is well defined from $C_{1}$ to $C_{0}$, it is an isomorphism and satisfies $\eta_{0}=h\circ\eta_{1}$.
      Similarly, note that the map $f_{0}:C_{0}\to E$ is the identity on $E$ and maps $\mathbb{P}^{1}$ to $P_{0}$, similarly the map $f_{1}:C_{1}\to E$ is the identity on $E$ and maps $\mathbb{P}^{1}$ to $P_{1}$. Therefore, we have the following commutative diagram, in which $h$ and $h_{a}$ are isomorphisms:
[…]
      It follows that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $\overline{M}_{E,g}/{\rm Aut}(E)]$. Since $\sigma_{a}^{*}$ and $\sigma_{b}^{*}$ determine $\eta_{0}^{*}(K(E))$ and $\eta_{0}^{*}(K(\mathbb{P}^{1}))$, we see that $D$ has only finitely many preimages.
    2. (b)
      If $g_{1}\geq 2$ and $g_{2}=1$. Then $(D^{\prime}\to C_{0})$ must be in $\Delta_{g-1,1}$ or $\Xi_{g-2,1}$. For now we consider both cases simultaneously and we distinguish between the different cases at the end. Note that there are only finitely many options for $\sigma_{a}$. Moreover, $D_{b}$ is an elliptic curve, so its $j$-invariant is the $j$-invariant of the cross-ratio of the four branch points of $\eta_{0}:D_{b}\to\mathbb{P}^{1}$. Thus there are at most six choices for this cross-ratio. We show that $\sigma_{a}$ and the cross ratio uniquely determine $D^{\prime}\to C_{0}\to E$.
      Let $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ be another preimage of $D$, and suppose that $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$ and $\eta_{0}:D_{b}\to\mathbb{P}^{1}$ and $\eta_{1}:D_{b}\to\mathbb{P}^{1}$ have the same cross-ratio of branch points. Then, similar to case (a), there is $h_{a}\in{\rm Aut}(E)$ such that $\eta_{1}^{*}\circ h_{a}^{*}=\eta_{0}^{*}$, so $h_{a}\circ\eta_{1}=\eta_{0}|_{D_{a}}$.
      Suppose $\eta_{0}:D_{b}\to\mathbb{P}^{1}$ is branched at $\{Q_{0},R_{0,1},R_{0,2},R_{0,3}\}$ and $\eta_{1}:D_{b}\to\mathbb{P}^{1}$ is branched at $\{Q_{1},R_{1,1},R_{1,2},R_{1,3}\}$. Since they have the same cross-ratio, there is a linear fractional transformation $h_{b}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ that sends $\{Q_{1},R_{1,1},R_{1,2},R_{1,3}\}$ to $\{Q_{0},R_{0,1},R_{0,2},R_{0,3}\}$. Now, $\eta_{0}$ and $h_{b}\circ\eta_{1}:D_{b}\to\mathbb{P}^{1}$ have the same branch points, so $\eta_{0}=h_{b}\circ\eta_{1}$. The rest of the argument is identical to case (a), which shows that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $\overline{M}_{E,g}/{\rm Aut}(E)]$.
      Now, when DD is in $\Delta_{g-1,1}$, then the point on $D_{b}$ that is clutched with $D_{a}$ is one of the four branch points $\{Q,R_{1},R_{2},R_{3}\}$, so there are finitely many possibilities. On the other hand, when DD is in $\Xi_{g-2,1}$, the clutching point of $D_{b}$ is any other point. The position of this point relative to $\{Q,R_{1},R_{2},R_{3}\}=D_{b}[2]$ introduces another dimension to the moduli problem. Two choices of a clutching point give the same point in $\overline{\mathcal{M}}_{g}$ if and only if there is an automorphism of $D_{b}$ that exchanges the two clutching points and preserves $D_{b}[2]$, so again this gives a finite number of preimages in $\overline{M}_{E,g}/{\rm Aut}(E)]$.
    3. (c)
      If $g_{1}\geq 2$ and $g_{2}=0$. Then, $(D^{\prime}\to C_{0})$ is in $\Xi_{g_{1},0}$, and DD is $D_{a}$ with two points in $\eta_{0}^{-1}(P_{0})$ identified with each other. Moreover, $g(D_{b})=0$, so $D_{b}=\mathbb{P}^{1}$. Again $\sigma_{a}$ has only finitely many choices.
      Suppose that $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage of $D$ for which $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$. Similar to case (a), there is $h_{a}\in{\rm Aut}(E)$ for which $h_{a}\circ\eta_{1}=\eta_{0}|_{D_{a}}$. This forces $h_{a}(P_{1})=P_{0}$. On the other hand, both $\eta_{0}|_{D_{b}}$ and $\eta_{1}|_{D_{b}}$ are double covers of $\mathbb{P}^{1}$ ramified at two points, so there is $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{0}=h_{b}\circ\eta_{1}$ and $h_{b}(Q_{1})=Q_{0}$. The rest of the proof is the same as in case (a) to show that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $\overline{M}_{E,g}/{\rm Aut}(E)]$.
    4. (d)
      If $g_{1}=1$ and $g_{2}\geq 2$. Then $\sigma_{b}$ has only finitely many choices and $\eta_{0}:D_{a}\to E$ has only $4$ choices (it is in $\overline{M}_{E,1}/{\rm Aut}(E)]$). Suppose that $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage for which $\eta_{0}^{*}(K(\mathbb{P}^{1}))=\eta_{1}^{*}(K(\mathbb{P}^{1}))$ and $\eta_{0},\eta_{1}:D_{a}\to E$ correspond to same point of $\overline{M}_{E,1}/{\rm Aut}(E)]$. Thus, there is $h_{a}\in{\rm Aut}(E)$ and $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{0}=h_{a}\circ\eta_{1}:D_{a}\to E$ and $\eta_{0}=h_{b}\circ\eta_{1}:D_{b}\to\mathbb{P}^{1}$. The rest of the proof is identical to case (a).
    5. (e)
      If $g_{1}=1$ and $g_{2}=1$, so $D^{\prime}\to E$ is in $\Delta_{1,1}$. This case is identical to case (b), except that $\eta_{0}:D_{a}\to E$ has only $4$ choices as an element of $\overline{M}_{E,1}/{\rm Aut}(E)]$.
    6. (f)
      If $g_{1}=1$ and $g_{2}=0$, so $D^{\prime}\to E$ is in $\Xi_{1,0}$. This case is identical to case (c), except that $\eta_{0}:D_{a}\to E$ has only $4$ choices as an element of $\overline{M}_{E,1}/{\rm Aut}(E)]$.
In all cases, we see that $D$ has only finitely many preimages in $\overline{M}_{E,g}/{\rm Aut}(E)]$. Thus, the map from $\overline{M}_{E,g}/{\rm Aut}(E)]$ to $\overline{\mathcal{M}}_{g}$ is generically quasi-finite on each boundary divisor. ∎